OpenStudy (celinegirl):

Is the simplified form of 2square root of 3 + 3square root of 3 rational? Yes No

OpenStudy (celinegirl):

I say no but I want to make sure im right

OpenStudy (anonymous):

\[2\sqrt{3}+3\sqrt{3}\]\[=5\sqrt{3}\]

OpenStudy (anonymous):

The square root of 3 times by 5 is irrational, so no

OpenStudy (celinegirl):

@livias.random Okay thank you